OpenStudy (anonymous):

integrate: sin3(2x)cos(2x)dx from 0 to 8

OpenStudy (anonymous):

like this? \[\int\limits_{0}^{8} \sin^3 (2x)\cos(2x)dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u-sub.. let u=sin(2x)

OpenStudy (anonymous):

i think i am doing it wrong. but i got 1/32

OpenStudy (anonymous):

opps and i mean pi/8...not 8

OpenStudy (anonymous):

pi/8 as the upper limit... integral = 1/32 is what i got too..

OpenStudy (anonymous):

sorry, but yes pi/8 is the upper limit. ok, thanks i'm just making sure i did it correctly....THANK YOU

OpenStudy (anonymous):

yw )